OpenStudy (anonymous):

solve -\frac{ 1 }{ 6 }m =22 A. 132 B. 22/6 C. -132 D. -22/6

OpenStudy (anonymous):

\[ -\frac{ 1 }{ 6 }m =22\]

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

(-1/6)m = 22 m = 22*(-6) m = ???

OpenStudy (anonymous):

-132